# On computability of equilibrium states

### v2

@@ 3.3. Computability of probability measures
Following [HR09, Proposition 4.1.3], if $(X,\,\rho,\,\mathcal{S})$ is a computable metric space, and $X$ is bounded, then $\bigl{(}\mathcal{P}(X),\,W_{\rho},\,\mathcal{R}_{\mathcal{S}}\bigr{)}$ is a computable metric space, where the Wasserstein–Kantorovich metric $W_{\rho}$ is recalled in (2.2), and $\mathcal{R}_{\mathcal{S}}\subseteq\mathcal{P}(X)$ is the set of the probability measures such that each measure is supported on finitely many points of $\mathcal{S}$ and assigns rational values to them.
@@
By [GHR10, Lemma 2.12 and Proposition 4], due to the completeness of $\mathcal{P}(X)$ with respect to the metric $W_{\rho}$, one can conclude the following result.
@@
Assume, in addition, that $X$ is recursively compact, then by [BRY12, Proposition 2.13], upper bounds of computable functions on $X$ can be computed. Hence, we have the following result.
@@
Finally, we introduce a family of computable functions.
@@
[…]
The hat function $g_{u,r,\epsilon}(x)$ is an $\epsilon^{-1}$-Lipschitz function which equals to $1$ in the ball $B(u,r)$, to 00 outside $B(u,r+\epsilon)$, and lies strictly between 00 and $1$ in $B(u,r+\epsilon)\setminus\overline{B(u,r)}$.
@@
[…]
According to the canonical bijection between $N^{*}$ and $N$ (see Subsection 2.1), it follows from Definitions 3.21 and 3.22 that there exists an enumeration $\{\varphi_{j}\}_{j\in\mathbb{N}}$ for $\mathcal{E}(\mathcal{S})$ such that $\{\varphi_{j}\}_{j\in\mathbb{N}}$ is a uniformly computable sequence of functions. We call such an enumeration a computable enumeration. Assume that $X$ is a compact separable metric space and $\mathcal{S}$ a dense subset of $X$, then it follows from the lattice version of the Stone–Weierstrass theorem that $\mathcal{E}(\mathcal{S})$ is dense in $C(X)$. Thus, using the dominated convergence theorem, we conclude the following proposition.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, $X$ be recursively compact, and $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be a computable enumeration of $\mathcal{E}(\mathcal{S})$. Then for each pair of Borel probability measures $\mu,\,\nu\in\mathcal{P}(X)$, $\mu=\nu$ if and only if $\int\!\varphi_{j}\,\mathrm{d}\mu=\int\!\varphi_{j}\,\mathrm{d}\nu$, for each $j\in\mathbb{N}.$
@@ 4. Computable Analysis on thermodynamic formalism
In this section, we establish Theorems 1.1 and 1.2. In Subsection 4.1, we estimate the rates of convergences of iterations of some normalized Ruelle–Perron–Frobenius operators. In Subsection 4.2, we prove that the Jacobian function $J_{\varphi}(x)$ (see (4.19)) is computable. In Subsection 4.3, we demonstrate that two subsets $U$ and $W$ of $\mathcal{P}(X)$ (from Lemmas 4.12 and 4.13) are both lower-computable open sets. In Subsection 4.4, we complete the proofs of Theorems 1.1 and 1.2.
@@ 4.1. Ruelle–Perron–Frobenius operators and cones
In this subsection, we develop the technique of cones, combining ideas from [PU10], to prove in Theorem 4.1 that the convergence of iterations of some normalized Ruelle–Perron–Frobenius operator is at an exponential rate with explicit expressions for the related constants. The technique of cones first used by Birkhoff [Bi57] in the study of Banach spaces geometry was reintroduced by Liverani [Liv95a, Liv95b] for the analysis of the decay of correlations. For a quick introduction to cones in the context of connected spaces, see [Vi97, Section 2.2]. The lack of connectedness of $X$ becomes a key obstacle in our study of Computable Analysis. Due to this obstacle and the demand on the computability (in the sense of Computable Analysis) of various constants involved, we need to further refine the technique of cones. In this subsection, we denote
(4.1) $\widehat{K}(\lambda_{1},b,b^{\prime})\coloneqq 2\log\biggl{(}\frac{1+\lambda_{1}}{1-\lambda_{1}}\cdot\frac{b^{2}}{b-b^{\prime}}\biggr{)}>0$
for $1\leqslant b^{\prime}<b$ and $\lambda_{1}\in(0,1)$.
@@
[…]
holds for each $k\in\mathbb{N}$ satisfying $0\leqslant Z(1-\exp(-Z))^{k-1}\leqslant 1$, where $u_{\varphi}$ is the function from Corollary 2.9 and $Z\coloneqq\widehat{K}\bigl{(}\lambda_{1},C^{\prime},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\bigr{)}>0$.
The proof of Theorem 4.1 will be given at the end of this subsection. We begin by introducing some notations in the cone technique. Let $E$ be a vector space over $\mathbb{R}$. A convex cone in $E$ is a subset $\mathcal{C}\subseteq E\setminus\{0\}$ satisfying the following properties:
@@
    $\overline{\mathcal{C}}\cap(-\overline{\mathcal{C}})=\{0\}$.
Here $\overline{\mathcal{C}}$ is the set consisting of $w\in E$ satisfying that there exist $v\in\mathcal{C}$ and $\{t_{n}\}_{n\in\mathbb{N}}\subseteq\mathbb{R}^{+}$ such that $w+t_{n}v\in\mathcal{C}$ for each $n\in\mathbb{N}$ and that $t_{n}$ converges to 00 as $n$ tends to +$\infty$+ ∞. Denote $-\overline{\mathcal{C}}\coloneqq\{-v\in E:v\in\overline{\mathcal{C}}\}$. Define
(4.2) $\alpha(v_{1},v_{2})\coloneqq\sup\{t>0:v_{2}-tv_{1}\in\mathcal{C}\}\quad\text{and}\quad\beta(v_{1},v_{2})\coloneqq\inf\{s>0:sv_{1}-v_{2}\in\mathcal{C}\}$
for all $v_{1},\,v_{2}\in\mathcal{C}$, with the conventions $\sup\emptyset=0,\,\inf\emptyset=+\infty$, and
(4.3) $\theta(v_{1},v_{2})\coloneqq\log(\beta(v_{1},v_{2}))-\log(\alpha(v_{1},v_{2})),$
with $\theta(v_{1},v_{2})=+\infty$ if $\alpha(v_{1},v_{2})=0$ or $\beta(v_{1},v_{2})=+\infty$. By the definition of a convex cone, we have $\alpha(v_{1},v_{2})\leqslant\beta(v_{1},v_{2}),\,\alpha(v_{1},v_{2})<+\infty,$ and $\beta(v_{1},v_{2})>0$ for all $v_{1},\,v_{2}\in\mathcal{C}$. Thus $\theta(v_{1},v_{2})$ is well-defined and takes value in $[0,+\infty]$. We call $\theta(\cdot,\cdot)$ the projective metric (on $\mathcal{C}$) associated to the convex cone $\mathcal{C}$.
@@
Let $\mathcal{C}$ be a convex cone. Then the projective metric $\theta$ has the following properties:
@@
    for all $v_{1},\,v_{2}\in\mathcal{C}$, $\theta(v_{1},v_{2})=0$ if and only if there exists $t>0$ such that $v_{1}=tv_{2}$.
[…]
@@
[…]
@@
To prove Theorem 4.1, we introduce the following family of convex cones.
@@
Let $(X,\rho)$ be a compact metric space. Given constants $\xi,\,a,\,v>0$, and $b>1$, denote by $\mathcal{C}(a,v,b)\subseteq C(X)$ the convex cone consisting of continuous functions $u\colon X\rightarrow\mathbb{R}$ satisfying the following properties:
@@
    $u(x)\leqslant\exp(a\rho(x,y)^{v})\cdot u(y)$ for all $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$,
@@
    $u(x)\leqslant bu(y)$ for each pair of $x,\,y\in X$,
and let $\alpha_{a,v,b}$ and $\beta_{a,v,b}$ be the corresponding functions associated to the convex cone $\mathcal{C}(a,v,b)$ defined in (4.2), and $\theta_{a,v,b}$ be the projective metric associated to the convex cone $\mathcal{C}(a,v,b)$ defined in (4.3).
Let us compute the expression of $\alpha_{a,v,b}$ for a given pair of $\varphi_{1},\varphi_{2}\in\mathcal{C}(a,v,b)$. By (4.2), $\alpha_{a,v,b}(\varphi_{1},\varphi_{2})$ is the supremum of all positive real numbers $t$ satisfying the following three properties for all $s,\,m,\,n,\,x,\,y\in X$ with $0<\rho(x,y)\leqslant\xi$:
@@
    $\frac{\varphi_{2}(x)-t\varphi_{1}(x)}{\varphi_{2}(y)-t\varphi_{1}(y)}\leqslant\exp(a\rho(x,y)^{v})$, or equivalently, $t\leqslant\frac{e^{a\rho(x,y)^{v}}\varphi_{2}(y)-\varphi_{2}(x)}{e^{a\rho(x,y)^{v}}\varphi_{1}(y)-\varphi_{1}(x)}$;
@@
In other words, $\alpha_{a,v,b}(\varphi_{1},\varphi_{2})$ is given by
(4.4) $\alpha_{a,v,b}(\varphi_{1},\varphi_{2})=\min\biggl{\{}\inf\limits_{x,\,y}\biggl{\{}\frac{e^{a\rho(x,y)^{v}}\varphi_{2}(x)-\varphi_{2}(y)}{e^{a\rho(x,y)^{v}}\varphi_{1}(x)-\varphi_{1}(y)}\biggr{\}},\,\inf\limits_{m,\,n\in X}\biggl{\{}\frac{b\varphi_{2}(m)-\varphi_{2}(n)}{b\varphi_{1}(m)-\varphi_{1}(n)}\biggr{\}},\,\inf\limits_{s\in X}\biggl{\{}\frac{\varphi_{2}(s)}{\varphi_{1}(s)}\biggr{\}}\biggr{\}},$
where the infimum in the first term is taken over all pairs of $x,\,y\in X$ with $0<\rho(x,y)\leqslant\xi$. We demonstrate that the images of the convex cone under iterations of the Ruelle–Perron–Frobenius operator contract uniformly.
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0},\,v_{0}>0$, and $\lambda>1$. Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ a continuous, topologically exact, open map which is distance-expanding with respect to the metric $\rho$ with constants $\eta$, $\lambda$, and $\xi$, and the potential $\varphi\colon X\rightarrow\mathbb{R}$ a Hölder continuous function with respect to the metric $\rho$ with exponent $v_{0}$ and constant $a_{0}$. Put $\lambda_{1}\coloneqq\frac{1+\lambda^{-v_{0}}}{2}<1$ and $a^{\prime}\coloneqq\frac{6a_{0}\lambda^{v_{0}}-2a_{0}}{(\lambda^{v_{0}}-1)^{2}}$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, and $Q\coloneqq\exp(\inf_{x\in X}\overline{\varphi}(x))>0$. Fix a constant $b>2\exp(a^{\prime})$. Suppose that $\epsilon\in(0,\xi]$ is a constant with $2\exp(a^{\prime}\epsilon^{v_{0}})\leqslant b$, and that $G^{\prime}=\{x^{\prime}_{1},\,x^{\prime}_{2},\,\ldots,\,x^{\prime}_{n^{\prime}}\}$ is an $(\epsilon/2)$-net of $X$. Fix an arbitrary integer $m\in\mathbb{N}$ satisfying that $\bigcap_{i=1}^{n^{\prime}}T^{m}(B(x^{\prime}_{i},\epsilon/2))=X.$ Put $b^{\prime}\coloneqq\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}<b$, where $C>1$ is from Lemma 2.8. Denote by $L\colon C(X)\rightarrow C(X)$ the operator given by
@@
for all $u\in C(X)$ and $x\in X$. Here $u_{\varphi}$ is the function from Corollary 2.9. Then $L(\mathbbm{1})=\mathbbm{1}$ and
(4.6) $L^{m}(\mathcal{C}(a^{\prime},v_{0},b))\subseteq\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})\subseteq\mathcal{C}(a^{\prime},v_{0},b).$
@@
By (2.15) and (4.5), we have $L(\mathbbm{1})=\mathbbm{1}$. It remains to prove (4.6). Since $\lambda_{1}<1$ and $b^{\prime}<b$, it follows from Definition 4.4 that $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})\subseteq\mathcal{C}(a^{\prime},v_{0},b)$. Fix an arbitrary $u\in\mathcal{C}(a^{\prime},v_{0},b)$. It suffices to demonstrate that $L^{m}(u)$ satisfies properties (i), (ii), and (iii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$. Since $u\in\mathcal{C}(a^{\prime},v_{0},b)$, then $u(x)>0$ for each $x\in X$. By (4.5), $L(u)(x)>0$ for each $x\in X$. By induction, we obtain that $L^{m}(u)(x)>0$ for each $x\in X$, so $L^{m}(u)$ satisfies property (i) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$. Next, we prove that $L(u)$ satisfies property (ii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$, namely,
@@
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. Consider a pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. Since $T$ is an open map which is distance-expanding with constants $\eta$, $\lambda$, and $\xi$, then for each $\overline{x}\in T^{-1}(x)$, one can find a point $y(\overline{x})\coloneqq T_{\overline{x}}^{-1}(y)\in T^{-1}(y)$ satisfying that $\rho(\overline{x},y(\overline{x}))\leqslant\lambda^{-1}\rho(x,y)\leqslant\xi,$ and hence,

### v3

@@ 3.3. Computability of probability measures
Following [HR09, Proposition 4.1.3], if $(X,\,\rho,\,\mathcal{S})$ is a computable metric space, and $X$ is bounded, then $\bigl{(}\mathcal{P}(X),\,W_{\rho},\,\mathcal{R}_{\mathcal{S}}\bigr{)}$ is a computable metric space, where the Wasserstein–Kantorovich metric $W_{\rho}$ is recalled in (2.1), and $\mathcal{R}_{\mathcal{S}}\subseteq\mathcal{P}(X)$ is the set of Borel probability measures $\mu$ satisfying $\operatorname{supp}(\mu)\subseteq\mathcal{S},\,\operatorname{card}(\operatorname{supp}(\mu))<+\infty$, and $\mu(\{x\})\in\mathbb{Q}$ for each $x\in\mathcal{S}$. Recall that the support $\operatorname{supp}(\mu)$ of a Borel measure $\mu$ is the complement of the union of all the open sets $U$ with $\mu(U)=0$.
@@
By [GHR11, Lemma 2.12 and Proposition 4], due to the completeness of $\mathcal{P}(X)$ with respect to the metric $W_{\rho}$, one can conclude the following result.
@@
Assume, in addition, that $X$ is recursively compact. Then by [BRY12, Proposition 2.13], upper bounds of computable functions on $X$ can be computed. Hence, we have the following result.
@@
Finally, we consider a family of computable functions.
@@
is called a hat function. Here $\lvert a\rvert^{+}\coloneqq\max\{a,\,0\}$.
The hat function $g_{u,r,\epsilon}(x)$ is an $\epsilon^{-1}$-Lipschitz continuous function that takes the value $1$ in the ball $B(u,r)$ and vanishes outside the enlarged ball $B(u,r+\epsilon)$.
@@
In a computable metric space $(X,\,\rho,\,\mathcal{S})$, assume that $\mathcal{F}_{0}(\mathcal{S})$ is the set of functions of the form $g_{u,\,r,\,1/n}(x)$, where $u\in\mathcal{S},\,r\in\mathbb{Q}^{+}$, $n\in\mathbb{N}$, and $\mathfrak{E}(\mathcal{S})$ is the smallest set of functions containing $\mathcal{F}_{0}$ and the constant function $1$ closed under $\max,\,\min$, and finite rational linear combinations. Then the elements in $\mathfrak{E}(\mathcal{S})$ are called test functions.
Remark.
Note that there exists a computable bijection between $\mathbb{N}^{*}$ and $\mathbb{N}$. Hence, from Definitions 3.21 and 3.22, it is not hard to construct an effective enumeration $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $\mathfrak{E}(\mathcal{S})$, i.e., $\{\varphi_{j}\}_{j\in\mathbb{N}}$ is a sequence of uniformly computable functions. We fix such an effective enumeration of $\mathfrak{E}(\mathcal{S})$ and call it the effective enumeration of $\mathfrak{E}(\mathcal{S})$ in $(X,\,\rho,\,\mathcal{S})$.
Assume that $X$ is a compact separable metric space and $\mathcal{S}$ a dense subset of $X$. Then by the Stone–Weierstrass theorem (see e.g. [Fo99, Theorem 4.45]), we have that $\mathfrak{E}(\mathcal{S})$ is dense in $C(X)$. Thus, using the dominated convergence theorem, we conclude the following proposition.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space, $X$ be recursively compact, and $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be a computable enumeration of $\mathfrak{E}(\mathcal{S})$. Then for each pair of Borel probability measures $\mu,\,\nu\in\mathcal{P}(X)$, $\mu=\nu$ if and only if $\int\!\varphi_{j}\,\mathrm{d}\mu=\int\!\varphi_{j}\,\mathrm{d}\nu$ for each $j\in\mathbb{N}.$
@@ 4. Computable Analysis on thermodynamic formalism
In this section, we develop the cone technique to establish Theorems 1.1 and 1.2. In Subsection 4.1, we estimate the rate of convergence of iterations of the normalized Ruelle–Perron–Frobenius operator. In Subsection 4.2, we prove that the Jacobian function $J_{\varphi}(x)$ (see (4.20)) is computable. In Subsection 4.3, we demonstrate that two subsets $U$ and $W$ of $\mathcal{P}(X)$ (from Lemmas 4.12 and 4.13) are both lower-computable open sets. In Subsection 4.4, we complete the proofs of Theorems 1.1 and 1.2.
@@ 4.1. Ruelle–Perron–Frobenius operators and cones
In this subsection, we develop the cone technique, combining ideas from [PU10], to prove in Theorem 4.1 that the convergence of iterations of the normalized Ruelle–Perron–Frobenius operator is at an exponential rate with explicit expression for the related constants. The cone technique, first used by Birkhoff [Bi57] in the study of Banach spaces geometry, was reintroduced by Liverani [Liv95a, Liv95b] for the research on the decay of correlations. For a quick introduction to the cone technique in the context of connected spaces, see [Vi97, Section 2.2]. The lack of connectedness of $X$ poses a key obstacle in our study of Computable Analysis. Due to this obstacle and the demand on the computability (in the sense of Computable Analysis) of various constants involved, we need to refine the cone technique further.
In this subsection, we denote
(4.1) $\widehat{K}(\lambda,b,b^{\prime})\coloneqq 2\log\biggl{(}\frac{1+\lambda}{1-\lambda}\cdot\frac{b^{2}}{b-b^{\prime}}\biggr{)}>0$
for all constants $\lambda,\,b,\,b^{\prime}\in\mathbb{R}$ with $1\leqslant b^{\prime}<b$ and $\lambda\in(0,1)$.
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0}>0,\,\lambda>1$, and $v_{0}\in(0,1]$. Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a continuous, topologically exact, open, and distance-expanding with constants $\eta$, $\lambda$, and $\xi$. Assume that $\varphi\in C^{0,v_{0}}(X,\rho)$ satisfies $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$.
Put $\lambda_{1}\coloneqq\frac{1+\lambda^{-v_{0}}}{2}<1$ and $a^{\prime}\coloneqq\frac{6a_{0}\lambda^{v_{0}}-2a_{0}}{(\lambda^{v_{0}}-1)^{2}}>0$. Fix an arbitrary constant $C^{\prime}>\max\bigl{\{}C^{2},\,2\exp(a^{\prime})\bigr{\}}$, where $C>1$ is the constant from Lemma 2.8. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, and $Q\coloneqq\exp(\inf_{x\in X}\overline{\varphi}(x))>0$. Suppose that $\epsilon\in(0,\xi)$ is a constant with $2\exp(a^{\prime}\epsilon^{v_{0}})\leqslant C^{\prime}$, and that $G^{\prime}=\{x^{\prime}_{i}:i\in[1,n^{\prime}]\cap\mathbb{N}\}$ is an $(\epsilon/2)$-net of $X$. Fix an arbitrary integer $m\in\mathbb{N}$ satisfying $\bigcap_{i=1}^{n^{\prime}}T^{m}(B(x^{\prime}_{i},\epsilon/2))=X.$
Then the following inequality
$\bigl{\|}\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})-u_{\varphi}\bigr{\|}_{\infty}\leqslant CeZ(1-\exp(-Z))^{k-1}$
holds for each $k\in\mathbb{N}$ with $0\leqslant Z(1-\exp(-Z))^{k-1}\leqslant 1$, where $u_{\varphi}$ is the function from Corollary 2.9 and $Z\coloneqq\widehat{K}\bigl{(}\lambda_{1},C^{\prime},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\bigr{)}>0$.
The proof of Theorem 4.1 will be given at the end of this subsection.
First, we introduce some notations in the cone technique. Let $E$ be a vector space over $\mathbb{R}$. A convex cone in $E$ is a subset $\mathcal{C}\subseteq E\smallsetminus\{0\}$ satisfying the following properties:
@@
    $\{v\in E:v\in\overline{\mathcal{C}},\,-v\in\overline{\mathcal{C}}\}=\{0\}$.
Here $\overline{\mathcal{C}}$ is the set consisting of $w\in E$ that satisfies that there exists $v\in\mathcal{C}$ and $\{t_{n}\}_{n\in\mathbb{N}}$ in $\mathbb{R}^{+}$ satisfying $w+t_{n}v\in\mathcal{C}$ for each $n\in\mathbb{N}$ and that $t_{n}$ converges to 0 as $n$ tends to $\infty$.
Consider an arbitrary pair of $v_{1},\,v_{2}\in\mathcal{C}$. Then we define
(4.2) $\alpha(v_{1},v_{2})\coloneqq\sup\{t>0:v_{2}-tv_{1}\in\mathcal{C}\}\quad\text{ and }\quad\beta(v_{1},v_{2})\coloneqq\inf\{s>0:sv_{1}-v_{2}\in\mathcal{C}\},$
with the conventions that $\sup\emptyset=0,\,\inf\emptyset=+\infty$. By the definition of convex cones, we have $\alpha(v_{1},v_{2})\leqslant\beta(v_{1},v_{2}),\,\alpha(v_{1},v_{2})<+\infty,$ and $\beta(v_{1},v_{2})>0$. Thus $\theta(v_{1},v_{2})$ given by
(4.3) $\theta(v_{1},v_{2})\coloneqq\begin{cases}+\infty&\text{if }\alpha(v_{1},v_{2})=0\text{ or }\beta(v_{1},v_{2})=+\infty;\\ \log(\beta(v_{1},v_{2}))-\log(\alpha(v_{1},v_{2}))&\text{otherwise}\end{cases}$
is well-defined and takes value in $0,+\infty]$. Here $\theta$ is called the projective metric associated to $\mathcal{C}$.
@@
Let $\mathcal{C}$ be a convex cone. Then the projective metric $\theta$ satisfies the following properties:
@@
    for all $v_{1},\,v_{2}\in\mathcal{C}$, $\theta(v_{1},v_{2})=0$ if and only if there exists $t>0$ with $v_{1}=tv_{2}$.
Moreover, the following proposition asserts that $L\colon\mathcal{C}_{1}\rightarrow\mathcal{C}_{2}$ is a strict contraction with respect to $\theta_{1}$ and $\theta_{2}$ if $L(\mathcal{C}_{1})$ has finite $\theta_{2}$-diameter. Here $\theta_{i}$ is the projective metric associated to $\mathcal{C}_{i}$ for each $i\in\{1,\,2\}$.
@@
Let $E_{i}$ be a vector space, and $\mathcal{C}_{i}\subseteq E_{i}$ be a convex cone with the projective metric $\theta_{i}$ for each $i\in\{1,\,2\}$. Assume that $L\colon E_{1}\rightarrow E_{2}$ is a linear operator with $L(\mathcal{C}_{1})\subseteq\mathcal{C}_{2}$, and $U\coloneqq\sup\{\theta_{2}(L(v),L(w)):v,\,w\in\mathcal{C}_{1}\}<+\infty$. Then for each pair of $v,\,w\in\mathcal{C}_{1}$,
@@
To prove Theorem 4.1, we introduce the following family of convex cones with the given constants $\xi>0$ and $v_{0}\in(0,1]$.
@@
Let $(X,\rho)$ be a compact metric space. For each $a>0$ and each $b>1$, denote by $\mathcal{C}(a,b)\subseteq C(X)$ the set consisting of continuous functions $u\colon X\rightarrow\mathbb{R}$ satisfying the following properties:
@@
    $u(x)\leqslant\exp(a\rho(x,y)^{v_{0}})\cdot u(y)$ for all $x,\,y\in X$ with $\rho(x,y)<\xi$,
@@
    $u(x)\leqslant bu(y)$ for each pair of $x,\,y\in X$.
It is not hard to show that $\mathcal{C}(a,b)$ is a convex cone. Let $\alpha_{a,b}$ and $\beta_{a,b}$ be the corresponding functions associated to $\mathcal{C}(a,b)$ defined in (4.2), and $\theta_{a,b}$ be the projective metric associated to $\mathcal{C}(a,b)$ defined in (4.3).
Consider an arbitrary pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(a,b).$ We shall compute the expression of $\alpha_{a,b}(\varphi_{1},\varphi_{2})$. By (4.2), $\alpha_{a,b}(\varphi_{1},\varphi_{2})$ is the supremum of all positive real numbers $t$ satisfying the following three properties for all $s,\,m,\,n,\,x,\,y\in X$ with $0<\rho(x,y)<\xi$:
@@
    $\frac{\varphi_{2}(x)-t\varphi_{1}(x)}{\varphi_{2}(y)-t\varphi_{1}(y)}\leqslant\exp(a\rho(x,y)^{v_{0}})$, or equivalently, $t\leqslant\frac{e^{a\rho(x,y)^{v_{0}}}\varphi_{2}(y)-\varphi_{2}(x)}{e^{a\rho(x,y)^{v_{0}}}\varphi_{1}(y)-\varphi_{1}(x)}$;
@@
In other words,
(4.4) $\alpha_{a,b}(\varphi_{1},\varphi_{2})=\min\biggl{\{}\inf\limits_{x,\,y}\biggl{\{}\frac{e^{a\rho(x,y)^{v_{0}}}\varphi_{2}(x)-\varphi_{2}(y)}{e^{a\rho(x,y)^{v_{0}}}\varphi_{1}(x)-\varphi_{1}(y)}\biggr{\}},\,\inf\limits_{m,\,n\in X}\biggl{\{}\frac{b\varphi_{2}(m)-\varphi_{2}(n)}{b\varphi_{1}(m)-\varphi_{1}(n)}\biggr{\}},\,\inf\limits_{s\in X}\biggl{\{}\frac{\varphi_{2}(s)}{\varphi_{1}(s)}\biggr{\}}\biggr{\}},$
where the infimum in the first term is taken over all pairs of $x,\,y\in X$ with $0<\rho(x,y)<\xi$.
Now we demonstrate that the images of a convex cone under iterations of the Ruelle–Perron–Frobenius operators contract uniformly.
@@
Fix arbitrary constants $\eta,\,\xi,\,a_{0}>0,\,\lambda>1$, and $v_{0}\in(0,1]$. Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a continuous, topologically exact, open, and distance-expanding with constants $\eta$, $\lambda$, and $\xi$. Assume that $\varphi\in C^{0,v_{0}}(X,\rho)$ satisfies $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$.
Put $\lambda_{1}\coloneqq\frac{1+\lambda^{-v_{0}}}{2}<1$ and $a^{\prime}\coloneqq\frac{6a_{0}\lambda^{v_{0}}-2a_{0}}{(\lambda^{v_{0}}-1)^{2}}>0$. Define $\overline{\varphi}(x)\coloneqq\varphi(x)-P(T,\varphi)$ for each $x\in X$, and $Q\coloneqq\exp(\inf_{x\in X}\overline{\varphi}(x))>0$. Fix a constant $b>2\exp(a^{\prime})$. Suppose that $\epsilon\in(0,\xi)$ is a constant with $2\exp(a^{\prime}\epsilon^{v_{0}})\leqslant b$, and that $G^{\prime}=\{x^{\prime}_{1},\,x^{\prime}_{2},\,\ldots,\,x^{\prime}_{n^{\prime}}\}$ is an $(\epsilon/2)$-net of $X$. Let $m\in\mathbb{N}$ be a constant satisfying $\bigcap_{i=1}^{n^{\prime}}T^{m}(B(x^{\prime}_{i},\epsilon/2))=X.$ Put $b^{\prime}\coloneqq\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}\in\bigl{(}\frac{b}{2},b\bigr{)}$, where $C>1$ is from Lemma 2.8.
Denote by $L\colon C(X)\rightarrow C(X)$ the operator given by
@@
for all $u\in C(X)$ and $x\in X$. Here $u_{\varphi}$ is the function from Corollary 2.9.
Then $L(\mathbbm{1})=\mathbbm{1}$ and
(4.6) $L^{m}(\mathcal{C}(a^{\prime},b))\subseteq\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})\subseteq\mathcal{C}(a^{\prime},b).$
@@
By (2.16) and (4.5), we obtain that $L(\mathbbm{1})=\mathbbm{1}$. It remains to prove (4.6). Since $\lambda_{1}<1$ and $b^{\prime}<b$, it follows from Definition 4.4 that $\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})\subseteq\mathcal{C}(a^{\prime},b)$. Hence, it suffices to show that $L^{m}(\mathcal{C}(a^{\prime},b))\subseteq\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})$.
Now we consider an arbitrary function $u\in\mathcal{C}(a^{\prime},b)$ and prove that $L^{m}(u)$ satisfies properties (i), (ii), and (iii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})$.
Since $u\in\mathcal{C}(a^{\prime},b)$, then $u(x)>0$ for each $x\in X$. By (4.5), $L(u)(x)>0$ for each $x\in X$. By induction, we obtain that $L^{m}(u)(x)>0$ for each $x\in X$, so $L^{m}(u)$ satisfies property (i) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})$.
Next, we prove that $L(u)$ satisfies property (ii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})$, namely,
@@
for each pair of $x,\,y\in X$ with $\rho(x,y)<\xi$. Consider a pair of $x,\,y\in X$ with $\rho(x,y)<\xi$. Since $T$ is an open and distance-expanding with constants $\eta$, $\lambda$, and $\xi$, for each $\overline{x}\in T^{-1}(x)$, there exists a point $y(\overline{x})\coloneqq T_{\overline{x}}^{-1}(y)\in T^{-1}(y)$ such that $\rho(\overline{x},y(\overline{x}))\leqslant\lambda^{-1}\rho(x,y)<\xi.$ Hence,